# Quantum heat engine in the optomechanical system with mechanical parametric drive

### v1

@@ III.2 Optimal Efficiency Control
Compared with the traditional thermodynamical cycle, a finite-time heat engine explicitly accounts for finite-time cycles, which generate non-zero power output. To obtain the higher output power, the evolution time in working processes is required as short as possible. In our model, due to the limit condition $\Omega_{t}>0$ to avoid the trap inversion [5], there exists the shortest evolution time in compression/expansion strokes. Below this limit, $\Omega_{t}$ becomes a complex value, the STA path breaks down.
@@
Fig. 5 illustrates the minimal evolution time $\tau_{\text{min}}$ in the working strokes. The corresponding efficiencies for the heat engine under the $\tau_{\text{min}}$ are plotted in Fig. 5, with the same corresponding efficiencies as in Fig. 4. It has been noticed that under the $\tau_{\text{min}}$ in the two working strokes, $\eta_{\text{STA}}<\eta_{\text{th}}$. For a time-dependent system, the faster time evolution usually has a larger non-adiabatic scattering rate, thus, the CD term needs higher extra energy to limit the system evolution close to the adiabatic eigenstates, the heat exchange induced by the STA path $\langle H^{c/e}_{\text{STA}}\rangle_{\tau}$, has a significant influence on the thermal efficiencies. Therefore, although the working processes have the largest output power, the STA path in $\tau_{\text{min}}$ leads to an efficiency even smaller than the non-adiabatic thermal efficiency.
Because the STA process leads to additional heat absorption in working strokes $\langle H^{c/e}_{\rm{STA}}\rangle_{\tau}$ is $\tau$-dependent. Therefore, a critical evolution time exists to distinguish whether the STA efficiency (Eq. (18)) is larger than the non-adiabatic thermal efficiency (Eq. (13)) or not. Under the critical evolution time $\tau_{\text{cri}}$, $\eta_{\text{th}}=\eta_{\text{STA}}$, which suggests
$\sum_{j=c,e}\langle H^{j}_{\text{STA}}\rangle_{\tau}=\frac{\langle W\rangle_{Q^{*}=1}\langle Q\rangle}{\langle W\rangle}-\langle Q\rangle_{Q^{*}=1}.$ (23)
The minimal $\tau$, which is obtained for the $\Omega_{t}$, is independent of the parametric drive strength $\lambda$. The critical time $\tau_{\text{cri}}$ is defined to hold the critical condition (Eq. (23)) that the contribution of heat absorption of the STA counteracts the non-adiabatic effect, which becomes larger when $\lambda$ increase, as plotted in Fig. 6. However, the optomechanical interaction has not been considered in the discussion for the $\tau_{\text{cri}}$ above. To consider the heat exchange $\langle\Delta E^{c/e}\rangle_{\tau}$ induced by the optomechanical interaction, because of $\bar{n}\gg\bar{n}_{c}$, in our system, the interaction leads to the extra heat absorption to the engine, with which $\sum_{j=c,e}\langle\Delta E^{j}\rangle_{\tau}>0$. Therefore, under the critical condition that ignored the optomechanical interaction, $\eta^{\prime}_{\text{STA}}>\eta^{\prime}_{\text{th}}$, as Fig. 6 suggested. It also suggests the optomechanical interaction decreases the $\tau_{\text{cri}}$.
Figure 6: (a). Minimal and critical evolution time. (b). Efficiencies at the critical time $\tau_{\text{cri}}$. The black dotted line is $\eta_{\text{th}}$; the green triangle is $\eta_{\text{STA}}$; the purple dashed line is the thermal efficiency within the optomechanical interaction, $\eta^{\prime}_{\text{th}}$; the blue circle is $\eta^{\prime}_{\text{STA}}$. Other parameters for both plots are the same as Fig. 4.
@@ Appendix A Heat flux
Here the $H_{\text{int}}$ and $H_{\text{PD}}$ correspond to the interaction and mechanical parametric drive terms, respectively. Each part of the above equations indicates the energy transformation between two subsystems $H_{a}$ and $H_{b}$, as illustrated in Fig. 1. If we define
@@ Appendix B The Fokker-Planck Equation
$\begin{split}\langle a^{\dagger}a\rangle_{ss}&=\frac{1}{2}\left.\big{\{}\frac{\kappa\bar{n}_{c}(\gamma-2\lambda)(\gamma+\kappa-2\lambda)+4G^{2}(\gamma\bar{n}+\kappa\bar{n}_{c}+\lambda)}{[4G^{2}+\kappa(\gamma-2\lambda)](\gamma+\kappa-2\lambda)}\right.\\ &+\left.\frac{\kappa\bar{n}_{c}(\gamma+2\lambda)(\gamma+\kappa+2\lambda)+4G^{2}(\gamma\bar{n}+\kappa\bar{n}_{c}-\lambda)}{[4G^{2}+\kappa(\gamma+2\lambda)](\gamma+\kappa+2\lambda)}\right.\big{\}},\end{split}$ (32)
$\begin{split}\langle b^{\dagger}b\rangle_{ss}&=\frac{1}{2}\left.\big{\{}\frac{\kappa(\gamma+\kappa-2\lambda)(\bar{n}\gamma+\lambda)+4G^{2}(\gamma\bar{n}+\kappa\bar{n}_{c}+\lambda)}{[4G^{2}+\kappa(\gamma-2\lambda)](\gamma+\kappa-2\lambda)}\right.\\ &+\left.\frac{\kappa(\gamma+\kappa+2\lambda)(\bar{n}\gamma-\lambda)+4G^{2}(\gamma\bar{n}+\kappa\bar{n}_{c}-\lambda)}{[4G^{2}+\kappa(\gamma+2\lambda)](\gamma+\kappa+2\lambda)}\right.\big{\}},\end{split}$ (33)
$\begin{split}\langle a^{\dagger}b\rangle_{ss}&=iG\kappa\left.\big{\{}\frac{(\bar{n}-\bar{n}_{c})\gamma+(1+2\bar{n}_{c})\lambda}{[4G^{2}+\kappa(\gamma-2\lambda)](\gamma+\kappa-2\lambda)}\right.\\ &+\left.\frac{(\bar{n}-\bar{n}_{c})\gamma-(1+2\bar{n}_{c})\lambda}{[4G^{2}+\kappa(\gamma+2\lambda)](\gamma+\kappa+2\lambda)}\right.\big{\}},\end{split}$ (34)
$\begin{split}\langle b^{2}\rangle_{ss}&=\left.\big{\{}\frac{\kappa(\gamma+\kappa-2\lambda)(\bar{n}\gamma+\lambda)+4G^{2}(\gamma\bar{n}+\kappa\bar{n}_{c}+\lambda)}{[4G^{2}+\kappa(\gamma-2\lambda)](\gamma+\kappa-2\lambda)}\right.\\ &-\left.\frac{\kappa(\gamma+\kappa+2\lambda)(\bar{n}\gamma-\lambda)+4G^{2}(\gamma\bar{n}+\kappa\bar{n}_{c}-\lambda)}{[4G^{2}+\kappa(\gamma+2\lambda)](\gamma+\kappa+2\lambda)}\right.\big{\}}.\end{split}$ (35)
@@
And $\vec{W}_{t}$ indicates n independent Wiener process, which in the short time limit is the term of Gaussian white noise. The matrix $A^{\text{int/0,c/e}}$,$B^{\text{c/e}}$ have the form

### v2

@@ III.2 Optimal Efficiency Control
Compared to the traditional thermodynamic cycle, a finite-time heat engine explicitly accounts for finite-time cycles, which generate a non-zero power output. To obtain the higher output power, the evolution time in working processes is required as short as possible. In our model, due to the limit condition $\Omega_{t}>0$ to avoid the trap inversion [21], there exists the shortest evolution time in compression/expansion strokes. Below this limit, $\Omega_{t}$ becomes a complex value, and the STA path breaks down.
@@
Fig. 5(a) illustrates the minimal evolution time $\tau_{\text{min}}$ in the working strokes. The corresponding efficiencies for the heat engine under the $\tau_{\text{min}}$ are plotted in Fig. 5(b), with the same corresponding efficiencies as in Fig. 4. It has been noticed that under the $\tau_{\text{min}}$ in the two working strokes, $\eta_{\text{STA}}<\eta_{\text{th}}$. For a time-dependent system, the faster time evolution usually has a larger non-adiabatic scattering rate, thus, the CD term needs higher extra energy to limit the system evolution close to the adiabatic eigenstates, the heat exchange induced by the STA path $\langle H^{c/e}_{\text{STA}}\rangle_{\tau}$, has a significant influence on the thermal efficiencies. Therefore, although the working processes have the largest output power, the STA path in $\tau_{\text{min}}$ leads to an efficiency even smaller than the non-adiabatic thermal efficiency.
Due to the $\tau$-dependency for the additional heat absorption $\langle H^{c/e}_{\rm{STA}}\rangle_{\tau}$ generated in working strokes, it can be deduced that there exists a critical evolution time $\tau_{\text{cri}}$ to distinguish whether the STA efficiency (Eq. (18)) is larger than the non-adiabatic thermal efficiency (Eq. (13)) or not, and for $\tau\lessgtr\tau_{\text{cri}}$, $\eta_{\text{th}}\gtrless\eta_{\text{SAT}}$. Under $\tau_{\text{cri}}$ that without optomechanical coupling, $\eta_{\text{th}}=\eta_{\text{STA}}$, which suggests
$\sum_{j=c,e}\langle H^{j}_{\text{STA}}\rangle_{\tau}=\frac{\langle W\rangle_{Q^{*}=1}\langle Q\rangle_{Q^{*}}}{\langle W\rangle_{Q^{*}}}-\langle Q\rangle_{Q^{*}=1}.$ (23)
The critical time $\tau_{\text{cri}}$ is defined to maintain the critical condition (Eq. (23)) that the contribution of heat absorption of the STA counteracts the non-adiabatic effect, which increases when $\lambda$ increases, as plotted in Fig. 6(a). Once the optomechanical coupling has been considered in the working strokes, the extra heat exchange $\langle\Delta E^{c/e}\rangle_{\tau}$ induced by the interaction leads to the extra heat absorption to the engine, with which $\sum_{j=c,e}\langle\Delta E^{j}\rangle_{\tau}>0$. This extra heat absorption requires more energy for the STA path to suppress the non-adiabatic scattering. Therefore, for the critical time $\tau^{\prime}_{\text{cri}}$ (corresponding to $\eta^{\prime}_{\text{STA}}=\eta^{\prime}_{\text{th}}$) when optomechanical coupling is taken into account, $\tau^{\prime}_{\text{cri}}<\tau_{\text{cri}}$, as illustrated in Fig. 6(b).
Figure 6: (a) Minimal and critical evolution time. (b) Efficiencies vs $\tau$. The dashed vertical lines indicate the critical time with ($\tau^{\prime}_{\text{cri}}$) and without ($\tau_{\text{cri}}$) optomechanical coupling. It can be noticed that the optomechanical interaction decreases the critical time. Other parameters are chosen the same as Fig. 4.
@@ Appendix A Heat flux
Here the $H_{\text{int}}$ and $H_{\text{PD}}$ correspond to the interaction and mechanical parametric drive terms, respectively. Each part of the above equations indicates the energy transformation between two subsystems $H_{a}$ and $H_{b}$, as illustrated in Fig. 1(a). If we define
@@ Appendix B The Fokker-Planck Equation
[…]
$\begin{split}\langle b^{2}\rangle_{ss}&=\left.\big{\{}\frac{\kappa(\gamma+\kappa-2\lambda)(\bar{n}\gamma+\lambda)+4G^{2}(\gamma\bar{n}+\kappa\bar{n}_{c}+\lambda)}{[4G^{2}+\kappa(\gamma-2\lambda)](\gamma+\kappa-2\lambda)}\right.\\ &-\left.\frac{\kappa(\gamma+\kappa+2\lambda)(\bar{n}\gamma-\lambda)+4G^{2}(\gamma\bar{n}+\kappa\bar{n}_{c}-\lambda)}{[4G^{2}+\kappa(\gamma+2\lambda)](\gamma+\kappa+2\lambda)}\right$$\big{\}}.\end{split}$ (35)
@@
And $\vec{W}_{t}$ indicates n independent Wiener process, which in the short time limit is the term of Gaussian white noise. The matrix $A^{\text{int/0,c/e}}$$B^{\text{c/e}}$ have the form